I have to solve the following question
It is a simultaneous equation using the method of substitution
The question is;
y^2 - x^2 = 60
and
y = 3x + 2
Please can you solve for X and Y!

Answer by jorel555(1290)

y=3x+2
(3x+2)²-x²-60=0
9x²+12x+4-x²-60=0
8x²+12x-56=0
2x²+3x-14=0
(2x+7)(x-2)=0
(x,y)=(-7/2,-17/2) or (2,8). ☺☺☺☺